Need ASAP!! Help please!!
4) Assume that you contribute an annual sum of money (your choice) to your companys retirement plan for the next 40 years and contribute that same sum of money to an identical investment outside of your companys retirement plan (non-tax deferred) for the same period of time, what would be the Future Value of each?
Answer #1

Answer:

A.

It is case of future value of annuity.

Taking annual savings = $2000 (my choice)

R = 3.4% ( choice)

Annual savings (P) = $2000

Time (n) = 40 years

Future value of money after 40 years = P*((1+R)^n – 1))/R = 2000*(1.034^40 – 1)/.034

Future value of money after 40 years = $165240.8